OpenStudy (anonymous):

Write an Equation for the line in slope intercept form y-0=8/9(x-0)

jimthompson5910 (jim_thompson5910):

\[\Large y - 0 = \frac{8}{9}(x-0)\] \[\Large y - 0 = \frac{8}{9}(x)+\frac{8}{9}(0)\] \[\Large y - 0 = \frac{8}{9}x\] \[\Large y = \frac{8}{9}x\]